# Image Watermarks are Removable Using Controllable Regeneration from Clean Noise

### v1

@@ Abstract
Image watermark techniques provide an effective way to assert ownership, deter misuse, and trace content sources, which has become increasingly essential in the era of large generative models. A critical attribute of watermark techniques is their robustness against various manipulations. In this paper, we introduce a watermark removal approach capable of effectively nullifying the state of the art watermarking techniques. Our primary insight involves regenerating the watermarked image starting from a clean Gaussian noise via a controllable diffusion model, utilizing the extracted semantic and spatial features from the watermarked image. The semantic control adapter and the spatial control network are specifically trained to control the denoising process towards ensuring image quality and enhancing consistency between the cleaned image and the original watermarked image. To achieve a smooth trade-off between watermark removal performance and image consistency, we further propose an adjustable and controllable regeneration scheme. This scheme adds varying numbers of noise steps to the latent representation of the watermarked image, followed by a controlled denoising process starting from this noisy latent representation. As the number of noise steps increases, the latent representation progressively approaches clean Gaussian noise, facilitating the desired trade-off. We apply our watermark removal methods across various watermarking techniques, and the results demonstrate that our methods offer superior visual consistency/quality and enhanced watermark removal performance compared to existing regeneration approaches.
@@ 1 Introduction
As the large generative models continue to advance, the realism of AI-generated content has reached unprecedented levels. Those AI-generated contents are nearly indistinguishable from content created by humans. While this technological progress brings about excitement and efficiency, the difficulty of distinguishing AI-generated and human-made content can lead to severe problems, such as the spread of misinformation and the potential erosion of trust in digital media. In response, watermarking AI-generated content (Wen et al., 2024; Zhao et al., 2023a; Saberi et al., 2023; Zhao et al., 2023b; Lukas et al., 2023; Yang et al., 2024; Ci et al., 2024b; a; Kirchenbauer et al., 2023; Zhang et al., 2024a; Liu & Bu, 2024; Rezaei et al., 2024) has emerged as an effective solution, providing a proactive and reliable method to embed hidden information into AI-generated content for identification and traceability.
@@ 3.3 Regeneration with control
Once our semantic control adapter and spatial control network are fully trained, they are ready to be deployed for direct watermark removal, requiring only a single watermarked image. The entire inference process of CtrlRegen is outlined in Algorithm 1. We sample the initial noise from a pure Gaussian distribution. Then the extracted image embedding and edge-detected image serve as conditions, providing the necessary semantic and spatial information. Two well-trained networks are integrated with the backbone SD to denoise across $T$ timesteps. Finally, the cleaned image is obtained by decoding the denoised latent representation.
The inference process of CtrlRegen+ is detailed in Algorithm 2. Unlike CtrlRegen, which utilizes pure Gaussian noise as a starting point, CtrlRegen+ first adds noise to the latent representation and then uses this noised representation as a starting point to reconstruct the image. The noising step can be selected based on the strength of different watermark methods. Algorithm 1 CtrlRegen 0:  Watermarked image $x_{w}$. 1:  $\hat{x}_{w}\leftarrow f(x_{w})$ 2:  $z_{T}\leftarrow\mathcal{N}(0,I_{n})$ 3:  for $t=T,T-1,\cdots,1$ do 4:     $\xi_{t}\leftarrow\epsilon_{\theta}(z_{t},x_{w},\hat{x}_{w},t)$ 5:     $z_{t-1}\leftarrow z_{t}-\xi_{t}$ 6:  end for 7:  $\tilde{x}\leftarrow\mathcal{D}(z_{0})$ 7:  Cleaned image $\tilde{x}$. Algorithm 2 CtrlRegen+ 0:  Watermarked image $x_{w}$, Noising step $t^{*}$. 1:  $\hat{x}_{w}\leftarrow f(x_{w})$, $z_{0}\leftarrow\mathcal{E}(x_{w})$ 2:  $\epsilon\leftarrow\mathcal{N}(0,I_{n})$, $z_{t^{*}}=\sqrt{\alpha_{t^{*}}}z_{0}+\sqrt{1-\alpha_{t^{*}}}\epsilon$ 3:  for $t=t^{*},\cdots,1$ do 4:     $\xi_{t}\leftarrow\epsilon_{\theta}(z_{t},x_{w},\hat{x}_{w},t)$ 5:     $z_{t-1}\leftarrow z_{t}-\xi_{t}$ 6:  end for 7:  $\tilde{x}\leftarrow\mathcal{D}(z_{0})$ 7:  Cleaned image $\tilde{x}$.
@@ 4.1 Experiment setting
Image Watermark Methods. To demonstrate the effectiveness of our watermark removal method, we evaluate it against seven diverse watermarking methods, including TreeRing (Wen et al., 2024), StableSignature (Fernandez et al., 2023), StegaStamp (Tancik et al., 2020), SSL (Fernandez et al., 2022), RivaGAN (Zhang et al., 2019) and DwtDctSvd (Cox et al., 2007). These methods encompass both low and high-perturbation watermarks, providing a comprehensive evaluation of our approach’s capabilities.
Image Watermark Removal Baselines. Our method is compared with two regeneration methods Regen (Zhao et al., 2023a) and Rinse (An et al., 2024). Regen employs the diffusion model to regenerate watermarked images through a process of noising and denoising. Rinse iteratively applies the Regen method multiple times to improve the efficacy of watermark removal. For Regen, we set the noising and denoising step as $70$, and Rinse implements the Regen twice.
@@ 4.2 Results
Visual Similarity and Quality of CtrlRegen. The regenerated images are assessed from two key aspects: visual similarity and quality. Table 1 presents these measurements for various watermark removal methods. For visual similarity, our method achieves a lower CLIP-FID score compared to Regen and Rinse. This lower score suggests that the distribution of our regenerated images more closely approximates that of the original watermarked images, reflecting better preservation of visual characteristics and overall image integrity. Our method exhibits a lower PSNR score for pixel-level measurement. However, it does not necessarily indicate severe degradation in similarity. The changes that lead to a lower PSNR can actually harmonize well with the content of the image without making the regenerated image appear unnatural. On the contrary, although the Regen and Rinse may achieve higher PSNR scores, it results in visible artifacts and degradation in the regenerated images compared to the original watermarked images, as illustrated in Figure 4. To further substantiate that images regenerated by our method exhibit superior image quality, we employ two image quality measurements: Q-Align and LIQE. These metrics are used to quantitatively assess and compare the quality of the regenerated images. From Table 1, our method achieves better image quality compared to Regen and Rinse.
@@
Adjustable and Controllable Regeneration. We evaluate our CtrlRegen+ against Regen across various aspects, including watermark removal effectiveness, visual resemblance, and image quality, by varying the number of noising steps. For high-perturbation watermarks, such as StegaStamp and TreeRing, we set the noising steps to be {$100$, $200$, $300$, $400$, $500$, $1000$} and sample from pure Gaussian noise. We do not set a noising step number larger than $500$ for Regen, as at this number of noising steps, the noised latent representation approaches pure Gaussian noise. Consequently, the uncontrolled regeneration process would likely produce an image completely unrelated to the original watermarked image.
@@ 5 Conclusion
In this paper, we introduce a more powerful regeneration attack that, combined with our proposed control techniques, effectively removes image watermarks by thoroughly eliminating the watermark information in the latent space. Unlike existing regeneration attacks that apply only a limited number of noising steps to the latent representation and thus can only be effective on low perturbation watermarks, our approach starts with clean noise. This method allows us to successfully invalidate even high-perturbation watermarks. Moreover, The introduced adjustable and controllable regeneration scheme allows attackers to tailor the attack strength according to the robustness of different watermarks. Our experiments demonstrate improved visual consistency and image quality compared to existing regeneration attacks under the same attack performance. Our attack is a no-box approach, meaning that the attacker only needs one watermarked image to remove watermarks without requiring knowledge of the watermarking scheme, detection rules, or access to the detector. By demonstrating the ability to defeat robust watermarking techniques, we highlight the urgent need for developing stronger watermarking solutions that can withstand these types of attacks.
Appendix A Appendix: Ablation Study
@@
Figure 7: Examples of CtrlRegen+ compared to Regen at different numbers of noising steps are illustrated, with the noising step $T$ set to {$100$, $200$, $300$, $400$, $500$}. The image is watermarked using StegaStamp and TreeRing.
Table 2: The visual similarity and quality performance between approaches using only semantic control versus those incorporating both semantic and spatial control.

### v2

@@ Abstract
Image watermark techniques provide an effective way to assert ownership, deter misuse, and trace content sources, which has become increasingly essential in the era of large generative models. A critical attribute of watermark techniques is their robustness against various manipulations. In this paper, we introduce a watermark removal approach capable of effectively nullifying state-of-the-art watermarking techniques. Our primary insight involves regenerating the watermarked image starting from a clean Gaussian noise via a controllable diffusion model, utilizing the extracted semantic and spatial features from the watermarked image. The semantic control adapter and the spatial control network are specifically trained to control the denoising process towards ensuring image quality and enhancing consistency between the cleaned image and the original watermarked image. To achieve a smooth trade-off between watermark removal performance and image consistency, we further propose an adjustable and controllable regeneration scheme. This scheme adds varying numbers of noise steps to the latent representation of the watermarked image, followed by a controlled denoising process starting from this noisy latent representation. As the number of noise steps increases, the latent representation progressively approaches clean Gaussian noise, facilitating the desired trade-off. We apply our watermark removal methods across various watermarking techniques, and the results demonstrate that our methods offer superior visual consistency/quality and enhanced watermark removal performance compared to existing regeneration approaches. Our code is available at https://github.com/yepengliu/CtrlRegen.
@@ 1 Introduction
As the large generative models continue to advance, the realism of AI-generated content has reached unprecedented levels. Those AI-generated contents are nearly indistinguishable from content created by humans. While this technological progress brings about excitement and efficiency, the difficulty of distinguishing AI-generated and human-made content can lead to severe problems, such as the spread of misinformation and the potential erosion of trust in digital media. In response, watermarking AI-generated content (Wen et al., 2024; Zhao et al., 2023a; Saberi et al., 2023; Zhao et al., 2023b; Lukas et al., 2023; Yang et al., 2024; Ci et al., 2024b; a; Kirchenbauer et al., 2023; Zhang et al., 2024a; Liu & Bu, 2024; Rezaei et al., 2024; He et al., 2024) has emerged as an effective solution, providing a proactive and reliable method to embed hidden information into AI-generated content for identification and traceability.
@@ 3.3 Regeneration with control
Once our semantic control adapter and spatial control network are fully trained, they are ready to be deployed for direct watermark removal, requiring only a single watermarked image. The entire inference process of CtrlRegen is outlined in Algorithm 1 of Appendix B. We sample the initial noise from a pure Gaussian distribution. Then the extracted image embedding and edge-detected image serve as conditions, providing the necessary semantic and spatial information. Two well-trained networks are integrated with the backbone SD to denoise across $T$ timesteps. Finally, the cleaned image is obtained by decoding the denoised latent representation.
The inference process of CtrlRegen+ is detailed in Algorithm 2 of Appendix B. Unlike CtrlRegen, which utilizes pure Gaussian noise as a starting point, CtrlRegen+ first adds noise to the latent representation and then uses this noised representation as a starting point to reconstruct the image. The noising step can be selected based on the strength of different watermark methods.
@@ 4.1 Experiment Setting
Image Watermark Methods. To demonstrate the effectiveness of our watermark removal method, we evaluate it against seven diverse watermarking methods, including TreeRing (Wen et al., 2024), StableSignature (Fernandez et al., 2023), StegaStamp (Tancik et al., 2020), HiDDeN (Zhu, 2018), SSL (Fernandez et al., 2022), RivaGAN (Zhang et al., 2019) and DwtDctSvd (Cox et al., 2007). These methods encompass both low and high-perturbation watermarks, providing a comprehensive evaluation of our approach’s capabilities. For multi-bit watermarking methods, the specific number of bits used for each method is provided in Table 4 of Appendix A.
Image Watermark Removal Baselines. Our method is compared with two regeneration methods Regen (Zhao et al., 2023a) and Rinse (An et al., 2024). Regen employs the diffusion model to regenerate watermarked images through a process of noising and denoising. Rinse iteratively applies the Regen method multiple times to improve the efficacy of watermark removal. For the experimental results shown in Table 1, we set the noising and denoising steps to 70 for Regen, while Rinse applies the Regen process twice.
@@ 4.2.1 CtrlRegen
Visual Similarity and Quality of CtrlRegen. The regenerated images are assessed from two key aspects: visual similarity and quality. Table 1 presents these measurements for various watermark removal methods. For visual similarity, our CtrlRegen achieves a lower CLIP-FID score compared to baselines. This lower score suggests that the distribution of our regenerated images more closely approximates that of the original watermarked images, reflecting better preservation of visual characteristics and overall image integrity. CtrlRegen exhibits a lower score for pixel-level measurement, PSNR, compared to baselines. However, it does not necessarily indicate severe degradation in similarity. The changes that lead to a lower PSNR can actually harmonize well with the content of the image without making the regenerated image appear unnatural. On the contrary, although the Regen and Rinse may achieve higher PSNR scores, it results in visible artifacts and degradation in the regenerated images compared to the original watermarked images, as illustrated in Figure 4. We provide more discussion and examples (Figure 8) for this problem in Appendix A. To further substantiate that images regenerated by our method exhibit superior image quality, we employ two image quality measurements: Q-Align and LIQE. These metrics are used to quantitatively assess and compare the quality of the regenerated images. From Table 1, our method achieves better image quality compared to Regen and Rinse.
@@
4.2.2 CtrlRegen+
We evaluate our CtrlRegen+ against Regen across various aspects, including watermark removal effectiveness, visual resemblance, and image quality, by varying the number of noising steps. For high-perturbation watermarks, such as StegaStamp and TreeRing, we set the noising steps to be {$100$, $200$, $300$, $400$, $500$, $1000$} and sample from pure Gaussian noise. We do not set a noising step number larger than $500$ for Regen, as at this number of noising steps, the noised latent representation approaches pure Gaussian noise. Consequently, the uncontrolled regeneration process would likely produce an image completely unrelated to the original watermarked image.
@@ 5 Conclusion
In this paper, we introduce a controllable regeneration attack that, combined with proposed control techniques, effectively removes image watermarks by thoroughly eliminating the watermark information in the latent space. Our experiments demonstrate improved visual consistency and image quality compared to existing uncontrolled regeneration attacks under the same removal performance. We note that our attack is a no-box approach, meaning that the attacker only needs one watermarked image to remove watermarks without requiring knowledge of the watermarking scheme, detection rules, or access to the detector. By demonstrating the ability to defeat robust watermarking techniques, we highlight the urgent need to develop stronger watermarking solutions that can withstand this attack. In this work, the consistency of regeneration may be constrained to a certain extent by the backbone model we employed. In future work, we will explore more advanced backbone models, such as SD-v3 (Esser et al., 2024), and refine our control techniques for further improvement.
Ethics Statement
Our primary motivation for developing a more effective watermark removal method is to uncover the vulnerabilities in current watermarking techniques. By demonstrating the weaknesses in existing watermarking methods, we aim to underscore the urgent need for more robust and resilient watermarking solutions. Additionally, our methods will serve as a benchmark for assessing the robustness of future watermarking techniques and fostering their advancement. The proposed watermark removal techniques may pose ethical risks, including potential copyright infringement, undermining digital rights management, and enabling the unauthorized use of protected content. It is essential that our methods should be used responsibly and in compliance with local regulations, ensuring this work contributes to strengthening, rather than compromising, digital watermarking security.
Acknowledgement
The UF team acknowledges UFIT Research Computing for providing computational resources and support that have contributed to the research results reported in this publication. The NUS team is only supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023).
Appendix A Additional Experimental Results and Ablation Study
@@
Figure 7: Visual examples of CtrlRegen+ compared to Regen at different numbers of noising steps are illustrated, with the noising step $T$ set to {$100$, $200$, $300$, $400$, $500$}. The image is watermarked using StegaStamp and TreeRing. It shows that our CtrlRegen+ preserves high visual consistency and quality, even across various noising steps.
@@
Table 2: The visual similarity and quality performance between approaches using only semantic control versus those incorporating both semantic and spatial control. The results reveal that semantic control alone provides only coarse-grained guidance, while incorporating both semantic and spatial controls significantly enhances image consistency and quality across various watermarking methods.
@@
Inference time of our watermark removal methods. To evaluate the inference time of our regeneration methods, we conducted experiments to compute the average inference time, including the backbone SD model using text as input, CtrlRegen, and CtrlRegen+ with different noising steps (200, 400, 600, 800, 1000). The results are presented in the following table. From Table 3, it is evident that our regeneration method (CtrlRegne, CtrlRegen+ 800, and CtrlRegen+ 1000) introduces a negligible time delay (less than 1 second) compared to the backbone Stable Diffusion model. Additionally, for CtrlRegen+ with noising steps smaller than 600, the inference time is even shorter than that of the backbone SD model. This demonstrates the efficiency of our approach.
Table 3: Inference time comparison of backbone SD, CtrlRegen, and CtrlRegen+ with varying noising steps for watermark removal. It highlights the efficiency of our methods.
Setting Backbone SD CtrlRegen CtrlRegen+ (200) CtrlRegen+ (400) CtrlRegen+ (600) CtrlRegen+ (800) CtrlRegen+ (1000)
Generation Time/s 1.74 2.56 0.69 1.19 1.71 2.15 2.64
Discussion about CtrlRegen Visual Similarity. In order to further illustrate that CtrlRegen preserves the overall content and ensures the perturbation integrates well with the image, we provide a visual comparison between CtrlRegen and Regen at the same PSNR level. Our experiments were conducted on StegaStamp watermarks, with Regen using a noising step of 360, while CtrlRegen generates images starting from clean noise. Under these settings, both methods achieve comparable PSNR values. However, as shown in Figure 8, at similar PSNR levels, Regen introduces noticeable distortions to the image content, whereas our method maintains strong alignment with the watermarked image. This comparison effectively demonstrates that the relatively lower PSNR of our approach does not compromise the overall integrity of the watermarked image content.
Figure 8: Visual examples of CtrlRegen compared to Regen at similar PSNR levels, where Regen uses a noising step of 360. The image is watermarked using StegaStamp. Regen introduces noticeable distortions to the image content, whereas our method maintains strong alignment with the watermarked image. This comparison effectively demonstrates that the relatively lower PSNR of our approach does not compromise the overall integrity of the watermarked image content.
Table 4: Bit number we used in our experiments for different watermarking methods.
Watermark DwtDctSvd RivaGAN SSL StableSignature StegaStamp
Bits 32 32 32 48 96
Appendix B Algorithm of proposed methods
Algorithm 1 CtrlRegen
0:  Watermarked image $x_{w}$.
1:  $\hat{x}_{w}\leftarrow f(x_{w})$
2:  $z_{T}\leftarrow\mathcal{N}(0,I_{n})$
3:  for $t=T,T-1,\cdots,1$ do
4:     $\xi_{t}\leftarrow\epsilon_{\theta}(z_{t},x_{w},\hat{x}_{w},t)$
5:     $z_{t-1}\leftarrow z_{t}-\xi_{t}$
6:  end for
7:  $\tilde{x}\leftarrow\mathcal{D}(z_{0})$
7:  Cleaned image $\tilde{x}$.
Algorithm 2 CtrlRegen+
0:  Watermarked image $x_{w}$, Noising step $t^{*}$.
1:  $\hat{x}_{w}\leftarrow f(x_{w})$, $z_{0}\leftarrow\mathcal{E}(x_{w})$
2:  $\epsilon\leftarrow\mathcal{N}(0,I_{n})$, $z_{t^{*}}=\sqrt{\alpha_{t^{*}}}z_{0}+\sqrt{1-\alpha_{t^{*}}}\epsilon$
3:  for $t=t^{*},\cdots,1$ do
4:     $\xi_{t}\leftarrow\epsilon_{\theta}(z_{t},x_{w},\hat{x}_{w},t)$
5:     $z_{t-1}\leftarrow z_{t}-\xi_{t}$
6:  end for
7:  $\tilde{x}\leftarrow\mathcal{D}(z_{0})$
7:  Cleaned image $\tilde{x}$.